OpenStudy (anonymous):

Express the greatest common divisor of the following pair of integers as a linear combination of the integers. 117, 213

OpenStudy (anonymous):

Your answer disappeared

OpenStudy (anonymous):

Sorry, I misinterpret the problem. We'll have to use the Euclidean algorithm.

OpenStudy (anonymous):

Okay Let me know what you get please.

OpenStudy (anonymous):

see the last example http://www.math.kent.edu/~soprunova/41001f11/notes_9_19.pdf

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

still stuck on 213 any help with that one?

OpenStudy (anonymous):

oldrin that example was good for 117 but not sure about 213

OpenStudy (anonymous):

$$213=117\cdot1+96\\117=96\cdot1+21\\96=21\cdot4+12\\21=12\cdot1+9\\12=9\cdot1+3\\9=3\cdot3+0$$... so \(3\) is our greatest common divisor. Now, we work backwards:$$96=213-117\\ 21=117-(213-117)=-213+2\cdot117\\ 12=(213-117)-4(-213+2\cdot117)=5\cdot213-9\cdot117\\ 9=(-213+2\cdot117)-(5\cdot213-9\cdot117)=-6\cdot213+11\cdot117\\ 3=(5\cdot213-9\cdot117)-(-6\cdot213+11\cdot117)=11\cdot213-20\cdot117$$

OpenStudy (anonymous):

Thank you so much